MW SPECTROSCOPY OF ALPHA-ALANINE AND A SEARCH IN ORI-KL

\underline{SHIORI WATANABE}, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; SATOSHI KUBOTA, KENTAROU KAWAGUCHI, Department of Chemistry, Faculty of Science, Okayama University, Tsushimanaka, Okayama 700-8530, Japan; YASUKO KA-SAI, SMILES Group, National Institute of Information and Communication Technology, Nukui-Kita, Koganei Tokyo, 184-8795, Japan; and TAKAMASA MOMOSE, Department of Chemistry, and Department of Physics and Astronomy, The University of British Columbia, Vancouver, B.C., V6T1Z1, Canada.

The study of amino acids in interstellar space is of great interest in connection with the origin of life. α-Alanine is the simplest amino acid that have a chiral carbon atom. Thus, the detection of α-alanine in interstellar space is essential to discuss the origin of the homo-chirality of natural amino acids in proteins. In order to search for α-alanine in interstellar space, we have observed rotational spectrum of α-alanine in the 83-99 GHz and 167-177 GHz regions by using a continuous molecular beam equipment. By combining previous data by Godfrey et al.^{a} and Blanco et al.^{b}, we obtained precise molecular constants that predict transitions in the 100 GHz region with an uncertainty of only 50 kHz. Based on the laboratory spectroscopy, we have searched for α-Alanine in Orion-KL by using the 45 m Nobeyama Radio Telescope. The results will be discussed.

aP. D. Godfrey, S. Firth, L. D. Hatherley, R. D. Brown, and A. P. Pierlot, J. Am. Chem. Soc. 115, 9687, (1993).

bS. Blanco, A. Lesarri, J. C. Lopez, and J. L. Alonso, J. Am. Chem. Soc. 126, 11675 (2004).